Find the inverse of the matrix, if it exists. A= -4 3 -5 1 s 8 5 3 32 32 3 32 8 5 32 8 5 32 3 32 1 32 Com
Answer #1

-4 3 A= -4 -5 d -b A- = ad-bc ad-bild -c a 3 ad-bca 64)65) - (3).(-4) - 20+19 = 32 -5 -3 Al 13 -u -5/32 -3/2 Hilar -4132 8 8